# SoK: Evaluating Jailbreak Guardrails for Large Language Models

### v1

@@ 4.2 Technical Paradigms
Rule-based Guardrails. These guardrails operate by employing predefined rules, patterns, or heuristics to detect potentially harmful inputs or outputs of LLMs. A typical rule-based approach includes utilizing keywords or regular expressions to identify specific patterns tied to harmful content. For instance, SmoothLLM, as referenced in [57], leverages keyword-based detectors to assess the safety of the LLM’s outputs, primarily to determine its jailbroken state. Similarly, the PII Detector mentioned in [60] uses regular expressions to identify personal identifiable information, such as phone numbers and emails. This method mirrors the approach taken by the baseline Regex in [78], which also utilizes regular expressions to mitigate jailbreak attacks.
Transitioning from specific examples to an evaluation of their effectiveness, it is evident that while these methods benefit from straightforward and transparent pattern matching—attributes that contribute to their computational efficiency and interpretability—their reliance on predefined patterns can be a significant drawback. Specifically, these rule-based systems may falter when encountering novel jailbreak techniques that deviate from known patterns which inherently limits their capability to combat more sophisticated attacks.
@@
Model-based Guardrails. These guardrails adopt classifiers or statistical characteristics to distinguish between benign and harmful queries. Model-based approaches can capture more complex patterns than rule-based methods, enabling them to generalize better to novel jailbreak attempts. Learning a text-based classifier is a common approach for jailbreak detection. On one hand, we can use traditional machine learning models as the classifiers [63, 66, 69, 76, 77, 78, 79]. For instance, K-Nearest Neighbors (KNN) in RigorLLM [63], LightGBM in RSAA [66] and Random Forest in PretrainedEmbeddings [79]. On the other hand, neural networks are also widely applied for the safety classification [51, 52, 56, 60, 65, 70, 76, 78, 87]. For example, HSF [76] and CURVALID [87] use a simple Multilayer Perceptron (MLP) as the classifier. PromptGuard [70] and erase-and-check [56] fine-tune the pre-trained model (i.e, mDeBERTa and DistilBERT, respectively) to distinguish the safe and unsafe inputs. Besides, other methods used statistical characteristics to design their own algorithms on safety distinguish [58, 61, 74, 75, 80, 85, 86, 88, 89]. Detecting Perplexity [54] and Perplexity Filter [55] classify the input as a jailbreak request if the perplexity is higher than a threshold. Gradient discrepancies between safe prompts and adversarial prompts are employed in GradSafe [58], GradientCuff [61] and TokenHighlighter [80]. JailGuard [89] identify the jailbroken state of responses by computing their KL-divergence. EEG-Defender [75], JBShield [85] and MirrorShield [88] take the model’s internal feature similarities between the input and the jailbreak prompt as judgment basis.
@@
Recent methods tend to conduct a detailed analysis before making a safety judgment. For example, SelfDefend [13] first summarizes the input intention and then assesses whether it constitutes a jailbreak request. GuardReasoner [83] and X-Guard [90] employ chain-of-thought reasoning to analyze the potential harms and finally give a safety judgment. More recently, however, there has been a discernible shift towards methodologies that conduct a more detailed, upfront analysis before arriving at a safety judgment. Illustrating this trend, SelfDefend [13] first summarizes the underlying intention of the input and then assesses whether this intention constitutes a jailbreak request. Building upon this principle of preliminary in-depth analysis, both GuardReasoner [83] and X-Guard [90] employ chain-of-thought reasoning. This allows them to meticulously trace and analyze potential harms associated with a query, culminating in a final safety judgment.
@@ 4.3 Safety Granularity
Session-level Guardrails. These guardrails monitor entire conversation sessions, tracking the evolution of dialogue across multiple turns to identify potential jailbreak attempts that unfold gradually. Session-level approaches can detect sophisticated multi-turn attacks that might appear benign when individual messages are analyzed in isolation. For instance, Adaptive Guardrail [74] maintain conversation state to identify harmful patterns across turns. These comprehensive approaches are particularly valuable against advanced jailbreak techniques that exploit the sequential nature of conversations but typically require more complex implementation and greater computational resources. We now present two RQs that explore the impact of safety granularity on the effectiveness and utility of guardrail mechanisms.
@@ 5.1 Evaluation Setup
Datasets and Target Models. Based on the five categories of existing jailbreak attacks we surveyed in §2 — manual, optimization-based, generation-based, implicit, and multi-turn jailbreaks — we identify representative jailbreak attack methods in each category. We then collect six benchmark datasets, JailbreakHub [20], JailbreakBench [92], SafeMTData [48], MultiJail [36], AlpacaEval [93] and OR-Bench [94], from which we use their user prompts for testing diverse guardrails. Table II lists the details of our collected benchmark datasets. JailbreakHub is a framework that collects and categorizes wild jailbreak prompts designed to bypass safety restrictions in LLMs. We randomly sample 1,000 in-the-wild prompts (IJP) from JailbreakHub as manual attacks. JailbreakBench is an open-source robustness benchmark specifically designed to evaluate and measure the vulnerability of LLMs to jailbreak attacks. We use a set of 100 harmful instructions from JailbreakBench to drive optimization-based jailbreaks (GCG [9] and AutoDAN [10]), generation-based jailbreaks (TAP [28] and LLM-Fuzzer [32]), implicit jailbreak attack DrAttack [34], and multi-turn jailbreak attack X-Teaming [49]. SafeMTData is a dataset contains the initial multi-turn jailbreak prompts produced by ActorAttack [48]. We choose 600 queries from SafeMTData as multi-turn jailbreak attacks. MultiJail is the first manually-created multilingual jailbreak dataset, covering both high-resource and low-resource languages. We choose 315 jailbreak prompts in Bengali from MultiJail as multilingual jailbreaks. AlpacaEval is an automatic evaluation framework designed to assess the performance of instruction-following LLMs. We use 805 instructions from AlpacaEval as normal prompts. OR-Bench is the first large-scale benchmark measuring LLMs’ over-refusal of 80,000 seemingly toxic but benign prompts across multiple categories. We randomly select 1,000 prompts from OR-Bench as normal prompts. Due to intra-processing guardrails are available to protect white-box LLMs, we evaluate the defense performance of guardrails on open-source LLMs, including Llama-3-8B-Instruct [6], and Vicuna-13b-v1.5 [4], covering diverse model architectures and sizes. Due to page limitation, readers may refer to Appendix Evaluation Results on Vicuna-13b-v1.5 for the results on Vicuna-13b-v1.5.
@@
Metrics. We measure the defense performance of guardrails by the attack success rate (ASR), i.e., the frequency with which jailbreak prompts in a benchmark dataset bypass the guardrail and the target LLM. The lower the ASR, the stronger the defense performance. Adopting the methodology from prior studies ([95, 96, 48, 49]), we leverage GPT-4o’s robust evaluation capabilities for assessment. This involves supplying GPT-4o, acting as a judge, with harmful objectives and the corresponding model outputs. The judge assigns a score from 1 to 5, where higher scores denote more harmful and detailed model responses. An attack is deemed successful if it receives a score of 5 from the GPT-4o Judge. For a detailed explanation of the scoring rubric, please see [95, 48].
@@
Attack Configuration. To assess the jailbreak defense performance of guardrails, we employ the most widely used jailbreak attacks, including a manual attack (IJP [20]), optimization-based attacks (GCG [9] and AutoDAN [10]), generation-based attacks (TAP [28] and LLM-Fuzzer [32]), implicit attacks (DrAttack [34] and MultiJail [36]), and multi-turn attacks (ActorAttack [48] and X-Teaming [49]). In the context of IJP, 1,000 adversarial queries were randomly sampled from the forbidden question set with jailbreak prompts [97], curated by JailbreakHub. Regarding GCG, its individual variant was selected, and the adversarial suffix was optimized against the target LLM employing a batch size of 512 and subjected to 500 optimization iterations. For the AutoDAN methodology, the hierarchically-guided genetic algorithm variant, specifically AutoDAN-HGA, was adopted. The genetic algorithm integral to AutoDAN-HGA operates with a crossover probability of 0.5, a mutation probability of 0.01, and undergoes 500 optimization iterations. Concerning TAP, the Vicuna-13b-v1.5 model [4] was utilized as the attacking agent. The parameters for TAP were configured with a maximum depth of 5, a maximum width of 5, and a branching factor of 4. The designated target models for TAP included Llama-3-8B-Instruct [6] or Vicuna-13b-v1.5 [4]. In the case of LLM-Fuzzer, GPT-3.5 served as the auxiliary model for generating mutational inputs, and the query limit directed at the target LLMs was established at 200. For DrAttack, jailbreak prompts were formulated using GPT-4o. With respect to MultiJail, the entirety of the 315 available queries in the Bengali language was selected. For the ActorAttack strategy, a corpus of 600 queries was sourced from the SafeMTData dataset [48] (specifically, the SafeMTData/Attack_600.json file available on Hugging Face). For X-Teaming, we set the attacking model as Qwen2.5-32B-Instruct [8] and use the TextGrad-based text optimization to refine jailbreak prompts. Regarding AlpacaEval, all 805 questions within the AlpacaEval dataset were utilized. For OR-Bench, a subset of 1,000 prompts was randomly selected from the OR-Bench dataset [94] (specifically, the or-bench-80k.csv file on Hugging Face).
@@
It is pertinent to note that: The prompts associated with IJP, MultiJail, ActorAttack, AlpacaEval, and OR-Bench are static in nature. Consequently, all guardrail mechanisms encounter identical input stimuli, irrespective of whether they are safeguarding Llama-3-8B or Vicuna-13b. In contrast, GCG, AutoDAN, and DrAttack are specifically tailored to either Llama-3-8B or Vicuna-13b. As such, guardrails receive uniform inputs when defending the same designated target LLM. Conversely, TAP, LLM-Fuzzer, and X-Teaming represent adaptive attack methodologies. This implies that guardrail systems are presented with varied inputs, even when applied to the identical target LLM.
Baselines. We compare our framework with popular jailbreak defense methods, including Perplexity Filter [55], SmoothLLM [57], Llama Guard [12], GradSafe [58], GradientCuff [61], SelfDefend [13], WildGuard [50], Prompt Guard [70], and GuardReasoner [83]. Specifically, Perplexity Filter leverages a Llama-2-7b model to calculate the perplexity of the input prompt. A jailbreak is considered to happen when the perplexity exceeds a threshold. We set this threshold at the maximum perplexity of any prompt in the JailbreakBench dataset of harmful behavior prompts. SmoothLLM perturbs the jailbreak prompts with character-level changes to enable the target LLM to perform defense. In this paper, we set SmoothLLM to conduct character swapping with a 10% perturbation percentage. Llama Guard is a fine-tuned Llama-2-7b model designed to detect the toxicity category of input prompts. GradSafe is a gradient-based detection method that identifies unsafe or jailbreak prompts in LLMs by analyzing the consistent gradient patterns of safety-critical parameters when paired with compliance responses. GradientCuff is a method for detecting jailbreak attacks on LLMs by analyzing the refusal loss landscape, leveraging gradient-based patterns to identify and block adversarial prompts while maintaining normal query performance. SelfDefend is a practical jailbreak defense framework for LLMs that uses a shadow LLM instance to concurrently detect harmful queries while the target LLM processes them, providing robust protection with minimal delay. WildGuard is an open, lightweight, multi-task moderation tool for LLMs that detects malicious user prompts, harmful model responses, and model refusal behavior. Prompt Guard is a security tool developed by Meta that detects and blocks malicious inputs (e.g., jailbreak attempts, prompt injections) in LLM applications, using lightweight classifier model Prompt-Guard-86M to filter harmful content in real time. GuardReasoner is a reasoning-based guard model designed to enhance the safety of LLMs by integrating explicit step-by-step reasoning into the moderation process. Our evaluations are implemented using PyTorch 2.6.0 and conducted on NVIDIA Hopper H800 GPUs.
@@
Figure 3: The heatmap of guardrails.

### v2

@@ 4.2 Technical Paradigms
Rule-based Guardrails. These guardrails operate by employing predefined rules, patterns, or heuristics to detect potentially harmful inputs or outputs of LLMs. A typical rule-based approach includes utilizing keywords or regular expressions to identify specific patterns tied to harmful content. For instance, SmoothLLM, as referenced in [67], leverages keyword-based detectors to assess the safety of the LLM’s outputs, primarily to determine its jailbroken state. Similarly, the PII Detector mentioned in [70] uses regular expressions to identify personally identifiable information, such as phone numbers and emails. This method mirrors the approach taken by the baseline Regex in [88], which also utilizes regular expressions to mitigate jailbreak attacks.
Transitioning from specific examples to an evaluation of their effectiveness, it is evident that while these methods benefit from straightforward and transparent pattern matching—attributes that contribute to their computational efficiency and interpretability—their reliance on predefined patterns can be a significant drawback. Specifically, these rule-based systems may falter when encountering novel jailbreak techniques that deviate from known patterns which inherently limit their capability to combat more sophisticated attacks.
@@
Summary 4: Rule-based guardrails, while beneficial for their computational efficiency and ease of interpretation, face challenges when dealing with innovative jailbreak techniques that do not match existing predefined patterns.
@@
Model-based Guardrails. These guardrails adopt classifiers or statistical characteristics to distinguish between benign and harmful queries. Model-based approaches can capture more complex patterns than rule-based methods, enabling them to generalize better to novel jailbreak attempts. Learning a text-based classifier is a common approach for jailbreak detection. On one hand, we can use traditional machine learning models as the classifiers [73, 76, 79, 86, 87, 88, 89]. For instance, K-Nearest Neighbors (KNN) in RigorLLM [73], LightGBM in RSAA [76] and Random Forest in PretrainedEmbeddings [89]. On the other hand, neural networks are also widely applied for the safety classification [61, 62, 66, 70, 75, 80, 86, 88, 96]. For example, HSF [86] and CURVALID [96] use a simple Multilayer Perceptron (MLP) as the classifier. PromptGuard [80] and erase-and-check [66] fine-tune the pre-trained model (i.e, mDeBERTa and DistilBERT, respectively) to distinguish the safe and unsafe inputs. Besides, other methods used statistical characteristics to design their own algorithms on safety distinguish [68, 71, 84, 85, 90, 13, 95, 97, 98]. Detecting Perplexity [64] and Perplexity Filter [65] classify the input as a jailbreak request if the perplexity is higher than a threshold. Gradient discrepancies between safe prompts and adversarial prompts are employed in GradSafe [68], GradientCuff [71] and TokenHighlighter [90]. JailGuard [98] identifies the jailbroken state of responses by computing their KL-divergence. EEG-Defender [85], JBShield [13] and MirrorShield [97] take the model’s internal feature similarities between the input and the jailbreak prompt as judgment basis.
@@
More recently, however, there has been a discernible shift towards methodologies that conduct a more detailed, upfront analysis before arriving at a safety judgment. Illustrating this trend, SelfDefend [11] first summarizes the underlying intention of the input and then assesses whether this intention constitutes a jailbreak request. Building upon this principle of preliminary in-depth analysis, both GuardReasoner [93] and X-Guard [99] employ chain-of-thought reasoning. This allows them to meticulously trace and analyze potential harms associated with a query, culminating in a final safety judgment.
@@ 4.3 Safety Granularity
Session-level Guardrails. These guardrails monitor entire conversation sessions, tracking the evolution of dialogue across multiple turns to identify potential jailbreak attempts that unfold gradually. Session-level approaches can detect sophisticated multi-turn attacks that might appear benign when individual messages are analyzed in isolation. For instance, Adaptive Guardrail [84] maintains conversation state to identify harmful patterns across turns. These comprehensive approaches are particularly valuable against advanced jailbreak techniques that exploit the sequential nature of conversations but typically require more complex implementation and greater computational resources. We now present two RQs that explore the impact of safety granularity on the effectiveness and utility of guardrail mechanisms.
@@
4.4 Reactiveness
Static Guardrails operate by analyzing inputs or outputs of LLMs without modifying them. For example, OpenAI Moderation [62] uses a fine-tuned model to classify input prompts into safety categories without altering the original query. In contrast, Dynamic Guardrails actively modify inputs or outputs to neutralize adversarial properties while preserving semantic meaning, which is beneficial for better identification of jailbreak attacks. The modification on the input prompt or the output response can be divided into character-level, token-level, word-level, and sentence-level changes.
Character-level changes involve randomly introducing and altering characters within the text input, such as typos, inserting new characters or character swaps, to disrupt adversarial prompts without changing the overall meaning [67, 69]. A prominent example is SmoothLLM [67], which perturbs multiple copies of the input prompt through character-level changes and uses a voting mechanism to determine safety. Token-level changes involve substituting, inserting, or deleting specific tokens to neutralize adversarial prompts. For instance, erase-and-check [66] removes potentially harmful tokens from the input prompt to construct multiple sanitized versions for safety voting like SmoothLLM. Word-level changes involve substituting words with synonyms or related terms to neutralize adversarial prompts [69, 98]. For instance, SemanticSmooth [69] converts all verbs into the past tense and substitutes both verbs & nouns with their semantic equivalents. Sentence-level changes involve modifying and rewriting the entire input query to expose the embedded attack intent [9, 69, 73, 88, 98, 99]. For example, NeMo Guardrails [9] encodes the input prompt into a structured format to clarify the request’s intent. RigorLLM [73] paraphrases or summarizes the input to identify potential jailbreak attempts. X-Guard [99] translates the non-English input into English. Naturally, textual modifications can be applied simultaneously at multiple levels of granularity, including character, word, and sentence levels, such as in SemanticSmooth [69] and JailGuard [98]. From a broader perspective, dynamic guardrails can be viewed as a form of input/output augmentation that enhances the robustness of jailbreak detection by exposing hidden attack intents.
4.5 Applicability
The applicability of a guardrail is largely determined by its dependency on model internals. White-box guardrails require access to the target LLM’s internal states, such as gradients or hidden activations, to detect jailbreak attempts. For instance, GradSafe [68] analyzes gradient patterns of safety-critical parameters, while JBShield [13] compares internal state divergences between benign and malicious queries. These methods can achieve high detection accuracy but are limited to scenarios where model internals are accessible. In contrast, black-box guardrails operate solely on input and output text, making them suitable for proprietary or API-based LLMs. Examples include Prompt Guard [80], which uses a lightweight classifier to filter malicious inputs, and WildGuard [60], which moderates both prompts and responses via a separate LLM. Black-box approaches offer broader deployment flexibility but may sacrifice granularity in detection. This dimension is critical for practical deployment: organizations using closed-source models like GPT-4 must rely on black-box guardrails, while those with custom LLMs can leverage white-box methods for finer control.
4.6 Explainability
Explainability refers to the ability of a guardrail to provide transparent, interpretable rationales for its safety judgments. Opaque guardrails, such as many model-based classifiers, output binary decisions without justification, which can hinder trust and debugging. For example, MoJE [87] employs simple linguistic statistical techniques to classify inputs as safe or unsafe without explaining its reasoning. While opaque guardrails can be efficient and effective, their lack of transparency may limit adoption in high-stakes applications where understanding failure modes is crucial. In contrast, explainable guardrails enhance usability by providing reasoning traces or confidence scores. GuardReasoner [93] employs chain-of-thought reasoning to step through potential harms before rendering a judgment, while SelfDefend [11] summarizes the intent of a query before assessing its risk. Such transparency helps developers understand failure modes, refine safety policies, and comply with regulatory requirements.
@@ 5.1 Evaluation Setup
Datasets and Target Models. Based on the five categories of existing jailbreak attacks we surveyed in §2 — manual, optimization-based, generation-based, implicit, and multi-turn jailbreaks — we identify representative jailbreak attack methods in each category. We then collect six benchmark datasets, JailbreakHub [17], JailbreakBench [101], SafeMTData [46], MultiJail [34], AlpacaEval [102] and OR-Bench [103], from which we use their user prompts to build test queries for diverse guardrails. Table II lists the details of our collected benchmark datasets. JailbreakHub is a framework that collects and categorizes wild jailbreak prompts designed to bypass safety restrictions in LLMs. To refine ASR/PGR to three decimal places, we randomly sample 1,000 in-the-wild prompts (IJP) from JailbreakHub as manual attacks. JailbreakBench is an open-source robustness benchmark specifically designed to evaluate and measure the vulnerability of LLMs to jailbreak attacks. We use all 100 harmful instructions from JailbreakBench to drive optimization-based jailbreaks (GCG [7] and AutoDAN [8]), generation-based jailbreaks (TAP [26] and LLM-Fuzzer [30]), implicit jailbreak attack DrAttack [32], and multi-turn jailbreak attack X-Teaming [47]. SafeMTData is a dataset contains the initial multi-turn jailbreak prompts produced by ActorAttack [46]. We choose all 600 queries from SafeMTData as multi-turn jailbreak attacks. MultiJail is the first manually-created multilingual jailbreak dataset, covering both high-resource and low-resource languages. We choose all 315 jailbreak prompts in Bengali from MultiJail as multilingual jailbreaks. AlpacaEval is an automatic evaluation framework designed to assess the performance of instruction-following LLMs. We use all 805 instructions from AlpacaEval as normal prompts. OR-Bench is the first large-scale benchmark measuring LLMs’ over-refusal of 80,000 seemingly toxic but benign prompts across multiple categories. We randomly select 1,000 prompts from OR-Bench as normal prompts to also calculate ASR/PGR to three decimal places. We evaluate the defense performance of guardrails on open-source LLMs (i.e., Llama-3-8B-Instruct [4] and Vicuna-13b-v1.5 [2]) and a closed-source LLM (i.e., GPT-4-0125-Preview [104]), covering diverse model architectures and sizes. For evaluations on Vicuna-13b-v1.5 and GPT-4-0125-Preview, please refer to Appendix A and Appendix B, respectively.
@@
Metrics. We measure the defense performance of guardrails by the attack success rate (ASR), i.e., the frequency with which jailbreak prompts in a benchmark dataset bypass the guardrail and the target LLM. The lower the ASR, the stronger the defense performance. Adopting the methodology from prior studies ([105, 46, 47]), we leverage GPT-4o’s robust evaluation capabilities for assessment. This involves supplying GPT-4o, acting as a judge, with harmful objectives and the corresponding model outputs. The judge assigns a score from 1 to 5, where higher scores denote more harmful and detailed model responses. An attack is deemed successful if it receives a score of 5 from the GPT-4o Judge. For a detailed explanation of the scoring rubric, please see [105, 46].
@@
Attack Configuration. To assess the jailbreak defense performance of guardrails, we employ the most widely used jailbreak attacks, including a manual attack (IJP [17]), optimization-based attacks (GCG [7] and AutoDAN [8]), generation-based attacks (TAP [26] and LLM-Fuzzer [30]), implicit attacks (DrAttack [32] and MultiJail [34]), and multi-turn attacks (ActorAttack [46] and X-Teaming [47]). In the context of IJP, 1,000 adversarial queries were randomly sampled from the forbidden question set with jailbreak prompts [106], curated by JailbreakHub. Regarding GCG, its individual variant was selected, and the adversarial suffix was optimized against the target LLM employing a batch size of 512 and subjected to 500 optimization iterations. For the AutoDAN methodology, the hierarchically-guided genetic algorithm variant, specifically AutoDAN-HGA, was adopted. The genetic algorithm integral to AutoDAN-HGA operates with a crossover probability of 0.5, a mutation probability of 0.01, and undergoes 500 optimization iterations. For GCG and AutoDAN, we migrate jailbreak prompts optimized for Vicuna-13b to attack GPT-4. Concerning TAP, the Vicuna-13b-v1.5 model [2] was utilized as the attacking agent. The parameters for TAP were configured with a maximum depth of 5, a maximum width of 5, and a branching factor of 4. The designated target models for TAP included Llama-3-8B-Instruct [4], Vicuna-13b-v1.5 [2] or GPT-4-0125-Preview [104]. In the case of LLM-Fuzzer, GPT-3.5 served as the auxiliary model for generating mutational inputs, and the query limit directed at the target LLMs was established at 200. For DrAttack, jailbreak prompts were formulated using GPT-4o. With respect to MultiJail, the entirety of the 315 available queries in the Bengali language was selected. For the ActorAttack strategy, a corpus of 600 queries was sourced from the SafeMTData dataset [46] (specifically, the SafeMTData/Attack_600.json file available on Hugging Face). For X-Teaming, we set the attacking model as Qwen2.5-32B-Instruct [6] and use the TextGrad-based text optimization to refine jailbreak prompts. Regarding AlpacaEval, all 805 questions within the AlpacaEval dataset were utilized. For OR-Bench, a subset of 1,000 prompts was randomly selected from the OR-Bench dataset [103] (specifically, the or-bench-80k.csv file on Hugging Face).
@@
It is pertinent to note that: The prompts associated with IJP, MultiJail, ActorAttack, AlpacaEval, and OR-Bench are static in nature. Consequently, all guardrail mechanisms encounter identical input stimuli, irrespective of whether they are safeguarding Llama-3-8B, Vicuna-13b or GPT-4. In contrast, GCG, AutoDAN, and DrAttack are specifically tailored to either Llama-3-8B, Vicuna-13b or GPT-4. As such, guardrails receive uniform inputs when defending the same designated target LLM. Conversely, TAP, LLM-Fuzzer, and X-Teaming represent adaptive attack methodologies. This implies that guardrail systems are presented with varied inputs, even when applied to the identical target LLM.
@@
Baselines. We evaluate our framework with popular jailbreak defense methods, including Perplexity Filter [65], SmoothLLM [67], Llama Guard [10], GradSafe [68], GradientCuff [71], SelfDefend [11], WildGuard [60], Prompt Guard [80], and GuardReasoner [93]. Specifically, Perplexity Filter leverages a Llama-2-7b model to calculate the perplexity of the input prompt. A jailbreak is considered to happen when the perplexity exceeds a threshold. We set this threshold at the maximum perplexity of any prompt in the JailbreakBench dataset of harmful behavior prompts. SmoothLLM perturbs the jailbreak prompts with character-level changes to enable the target LLM to perform defense. In this paper, we set SmoothLLM to conduct character swapping with a 10% perturbation percentage. Llama Guard is a fine-tuned Llama-2-7b model designed to detect the toxicity category of input prompts. GradSafe is a gradient-based detection method that identifies unsafe or jailbreak prompts in LLMs by analyzing the consistent gradient patterns of safety-critical parameters when paired with compliance responses. GradientCuff is a method for detecting jailbreak attacks on LLMs by analyzing the refusal loss landscape, leveraging gradient-based patterns to identify and block adversarial prompts while maintaining normal query performance. SelfDefend is a practical jailbreak defense framework for LLMs that uses a shadow LLM instance to concurrently detect harmful queries while the target LLM processes them, providing robust protection with minimal delay. WildGuard is an open, lightweight, multi-task moderation tool for LLMs that detects malicious user prompts, harmful model responses, and model refusal behavior. Prompt Guard is a security tool developed by Meta that detects and blocks malicious inputs (e.g., jailbreak attempts, prompt injections) in LLM applications, using a lightweight classifier model Prompt-Guard-86M to filter harmful content in real time. GuardReasoner is a reasoning-based guard model designed to enhance the safety of LLMs by integrating explicit step-by-step reasoning into the moderation process. Our evaluations are implemented using PyTorch 2.6.0 and conducted on NVIDIA Hopper H800 GPUs.